hi! how do u make a algebraic equation on this question: If the sides of a square are increased by 5cm, the area of the new square is four times the area of the original square. Find the length of a side of the original square.
Can you please show me how to solve this using the check. thanks!

Answer by user_dude2008(1862)

equation is (x+5)^2 = 4x^2

(x+5)^2 = 4x^2
x^2+10x+25 = 4x^2
-3x^2+10x+25 = 0
(x-5)(3x+5) = 0
x-5=0 or 3x+5=0
x=5 or x=-5/3

Only answer is x=5. Original side length is 5 cm

Check:
Original area = 5*5 = 25
New area = (5+5)*(5+5) = 10*10 = 100 = 4*25